### v2

# On-Device LLMs for Home Assistant: Dual Role in Intent Detection and Response Generation

Abstract
This paper investigates whether Large Language Models (LLMs), fine-tuned on synthetic but domain-representative data, can perform the twofold task of (i) slot and intent detection and (ii) natural language response generation for a smart home assistant, while running solely on resource-limited, CPU-only edge hardware. We fine-tune LLMs to produce both JSON action calls and text responses. Our experiments show that 16-bit and 8-bit quantized variants preserve high accuracy on slot and intent detection and maintain strong semantic coherence in generated text, while the 4-bit model, while retaining generative fluency, suffers a noticeable drop in device-service classification accuracy. Further evaluations on noisy human (non-synthetic) prompts and out-of-domain intents confirm the models’ generalization ability, obtaining around 80–86% accuracy. While the average inference time is 5–6 seconds per query—acceptable for one-shot commands but suboptimal for multi-turn dialogue—our results affirm that an on-device LLM can effectively unify command interpretation and flexible response generation for home automation without relying on specialized hardware.
On-Device LLMs for Home Assistant:
Dual Role in Intent Detection and Response Generation
Rune Birkmose*     Nathan Mørkeberg Reece*     Esben Hofstedt Norvin* Johannes Bjerva     Mike Zhang Aalborg University, Denmark *{rbirkm20, nreece20, enorvi20}@student.aau.dk   jjz@cs.aau.dk
**footnotetext: The authors contributed equally to this work.
1 Introduction
Smart home technologies and IoT devices have proliferated in recent years, with an expected rise from 16.6 billion to 18.8 billion connected devices by the end of 2024 (IoT Analytics, 2024). Major providers like Amazon, Google, and Apple typically handle speech recognition and intent detection on cloud servers, which raises user concerns about privacy, data ownership, and reliance on proprietary ecosystems (BBC News, 2025). Conventional solutions for home assistants often rely on specialized, domain-specific classifiers for slot and intent detection (SID), paired with templated system responses. While these approaches can be efficient, they can also be rigid, sometimes requiring precisely phrased user inputs and yielding repetitive or unpersonalized answers.
Recent developments in on-device computing—coupled with improvements in model compression and quantization Liang et al. (2021); Gholami et al. (2022); Lang et al. (2024)—have paved the way for smaller yet still capable language models to run on commodity hardware. These models offer privacy benefits and allow customizable local inference with reduced latency. However, deploying a capable model under strict memory and computational constraints remains challenging. Large-scale Transformer-based language models (Vaswani et al., 2017), and especially LLMs (Touvron et al., 2023; Dubey et al., 2024; Bai et al., 2023; Yang et al., 2024; Groeneveld et al., 2024), have demonstrated remarkable proficiency in tasks ranging from question answering to text generation (Arora et al., 2024; Yin et al., 2024), yet typically demand substantial hardware resources, restricting them to cloud-based services or large compute clusters.
This paper explores whether a smaller, fine-tuned LLM can provide two capabilities essential to a home assistant—accurate recognition of what users want (i.e., slot and intent detection), and natural textual responses—while running entirely on an edge device with limited CPU and memory. By unifying these tasks into one end-to-end system, we eliminate the need for separate domain-specific classification modules and templated responses, focusing on efficiency, robust language understanding, and strict correctness in JSON action output.
Additionally, we move away from classic SID datasets and other general spoken language understanding benchmarks. Instead, we investigate whether LLMs can be directly applied to digital assistant software. To this end, we take the open-source Home Assistant software [^1] https://github.com/home-assistant/core as our gold standard for evaluation, targeting real-world device-service pairs and actionable JSON outputs.
Contributions.
Our contributions can be summarized as follows: [^2] We release all our code and models at https://github.com/Run396/P9. 1 We show that a 0.5B LLM can be fine-tuned to already jointly handle SID and response generation with high accuracy. 2 By quantizing the model (from 16-bit to 8-bit and 4-bit), we quantify trade-offs between memory usage, accuracy, and generative fluency on CPU-only edge hardware. 3 We evaluate the approach on synthetic data, human queries, and out-of-domain tasks, confirming robust generalization.
2 Related Work
Slot and Intent Detection.
Traditional approaches to spoken language understanding (SLU) often treat SID separately using domain-specific classification or sequence tagging approaches (Zhang and Wang, 2016; Wang et al., 2018; Weld et al., 2022; Qin et al., 2021; Pham et al., 2023). More recent transformer-based solutions unify both tasks, leveraging contextual embeddings to improve performance (Castellucci et al., 2019; van der Goot et al., 2021; Stoica et al., 2021; Arora et al., 2024) with models like BERT (Devlin et al., 2019). However, many of these solutions still presume tailored sequence labeling datasets or full-size transformer backends. Our work aligns with the shift to more expressive transformer models for SLU, but we push inference to a local environment while also adding dynamic text generation.
Running LLMs on Edge Devices.
While training large-scale LLMs remains computationally expensive, numerous works explore strategies for deploying them on edge hardware. Haris et al. (2024) propose FPGA-based accelerators to reduce memory overhead for LLM inference. Zhang et al. (2024) distribute an LLM across multiple low-power devices to increase throughput. An empirical footprint study by Dhar et al. (2024) shows that even 7B-parameter models can strain embedded hardware if not sufficiently compressed. Our approach uses a much smaller LLM (0.5B–1.5B parameters) plus weight quantization, showing that near-commodity devices with 8GB RAM can handle both intent classification and text generation if the domain is sufficiently specialized.
3 Methodology
Our goal is to integrate two core functionalities of a home assistant into a single model:
  • Slot and Intent Detection: The model outputs a valid JSON object that maps to a desired service (intent) and device (slot) pair: {minted}[frame=single, framesep=2mm, xleftmargin=0pt, tabsize=2]js "service": "light.turn_on", "device": "light.living_room", "assistant": "Sure, turning on on the living room light."
  • Natural Language Generation: The model also produces a textual response confirming or elaborating on its action, as can be seen in the example above. The text can then be propagated to, e.g., a text-to-speech model.
Traditional classifiers only handle device-service classification and do not produce any text. For user-facing text, the baseline approach would rely on templated responses.
Partition Train Test Total
Classification 23,372 5,843 29,215
LLM 33,361 2,435 35,796
Table 1: Aggregated Train/Test Splits. For the classification baseline, 20% of the original training set was used as test data (after removing multi-intent samples). The LLM used the full synthetic data; 2,435 remain as test.
3.1 Data and Pre-processing
To the best of our knowledge, there is no existing human-curated dataset specifically for the Home Assistant software. Thus, we rely on synthetic data. We use a publicly available synthetic dataset (acon96, 2024), which consists of 35,840 synthetic examples designed to mimic Home Assistant commands. Each instance consists of:
  • A User Prompt: e.g., “Turn on the kitchen light”, “Set the thermostat to 22 degrees”.
  • One Valid JSON Action, containing the service and device fields corresponding to Home Assistant calls.
  • A Natural Language Response: e.g., a paraphrase or affirmation of the action taken.
A full example can be found in Figure 1 (Appendix A). We stratify the dataset, maintaining the inherent imbalance (some device types and services appear more frequently). There are 38 service labels and 858 device labels. We split into training and test sets as shown in Table 1. The final training set for the LLM includes $\sim$33k examples, and we set aside 2,435 synthetic samples for evaluation. Note that for the classification-based baselines, we split up the train and test set to separately predict service and device instead of as one prediction, ending up with double the test data (5,843 samples; excluding multi-intent examples). The input consists of only the user message and leave the system message out. A more detailed distribution of the data can be found in Table 6 (Appendix B).
Model Accuracy BERTScore
Baselines
SVC Classifier
      Service 76.6
      Device 45.4
DistilBERT
      Service 98.8
      Device 47.9
Qwen2.5-0.5B (16-bit) 98.8 0.84
Qwen2.5-0.5B (8-bit) 98.4 0.79
Qwen2.5-0.5B (4-bit) 81.7 0.88
Qwen2.5-1.5B (16-bit) 96.9 0.84
Qwen2.5-1.5B (8-bit) 96.5 0.83
Qwen2.5-1.5B (4-bit) 90.7 0.82
Table 2: Slot/Intent Detection and NLG Results on Synthetic Test Data. Accuracy is based on exact JSON match. BERTScore measures semantic similarity of the generated text vs. gold reference.
3.2 Models
Baseline Classifiers.
We train a Linear SVC from Scikit-Learn (Pedregosa et al., 2011) on TF-IDF features of the user prompt. The classifier outputs a concatenated device-service pair, which is then wrapped in JSON. Additionally, we fine-tune DistilBERT Sanh et al. (2019) for classification. We use the transformers library (Wolf et al., 2020) for fine-tuning. We train for 1 epoch using a learning rate of 3×$\times 10^{-4}$ with the AdamW optimizer, and a batch size of 64 on a NVIDIA A10 (24GB) GPUs. Both models have no generative capability, so user-facing text is templated.
Small Large Language Models.
We train using a chat-style format with user–assistant pairs. We primarily use the Qwen2.5-0.5B-Instruct model and the Qwen2.5-1.5B-Instruct (Yang et al., 2024). We fine-tune both models for one epoch with a batch size of 4, using the AdamW optimizer at a learning rate of $2\times 10^{-5}$ with a cosine scheduler. The maximum sequence length is set to 2,048 tokens. We use the HuggingFace Transformers library (Wolf et al., 2020) for training on NVIDIA L4 (24 GB) GPUs.
Quantization.
After fine-tuning and having the original 16-bit model, we produce two quantized versions of each model: NF8 and NF4 Dettmers et al. (2024), using bitsandbytes. [^3] See https://github.com/bitsandbytes-foundation/bitsandbytes. We also use double quantization. This allows us to compare accuracy, generative quality, and inference speed under varying memory constraints.
Model CPU T/Q (s) Load (s)
Baselines
SVC Classifier 4 $<$1
DistilBERT 4 $<$1
Qwen2.5-0.5B (16-bit) 4 6.25 $\pm$3.2
Qwen2.5-1.5B (16-bit) 4 10.81 $\pm$5.6
Qwen2.5-0.5B (8-bit) 4 5.50 $\pm$3.2
Qwen2.5-1.5B (8-bit) 4 10.32 $\pm$5.6
Qwen2.5-0.5B (16-bit) 2 8.49 $\pm$5.6
Qwen2.5-1.5B (16-bit) 2 17.72 $\pm$5.6
Qwen2.5-0.5B (8-bit) 2 7.89 $\pm$5.6
Qwen2.5-1.5B (8-bit) 2 16.11 $\pm$5.6
Table 3: Computation Time. Mean time per query (T/Q) across 500 samples under different CPU core counts and quantization levels. Load time is model initialization.
3.3 Evaluation
Slot-Intent Detection Accuracy.
SID must be correct with near-exact string matching, as JSON calls are consumed downstream by the home automation system. We thus parse the model output for the service and device fields; if they match the gold annotation exactly, it is counted as correct. Any mismatch or invalid JSON results in an error.
For the classification task, instead, we separately predict service and device using the same classification model and take the average accuracy.
Text Generation Quality.
For the natural language responses using the LLMs, we compare each generated response to the reference using BERTScore (Zhang et al., 2020).
Inference Environment.
We simulate a CPU-only setup on an 8 GB RAM device with up to four CPU cores. We measure average inference time on a 500-sample subset, varying both quantization level and the number of CPU cores.
Model Accuracy BERTScore
Qwen2.5-0.5B 80.0 0.76
Qwen2.5-1.5B 86.7 0.74
Table 4: Results Out-of-Domain Queries. Accuracy and BERTScore over 60 OOD samples.
4 Results
4.1 Slot and Intent Detection
Table 2 shows the SID performance of both the 0.5B and 1.5B LLMs under various quantization levels, alongside the baseline SVC and DistilBERT. For the 0.5B model, the 16-bit and 8-bit variants reach near-perfect accuracy ($\sim 99\%$). The 4-bit version drops to 81.7%, which is still better than the the SVC baseline (average 61.0% accuracy) and DistilBERT baseline (average 73.4% accuracy).
Interestingly, for the larger 1.5B model, the 16-bit and 8-bit variants achieve 96.9% and 96.5% accuracy, respectively, while the 4-bit version gets 90.7%. Thus, while the smaller 0.5B model actually yields higher raw accuracy in-domain, the 1.5B model remains competitive and in some out-of-domain tests (next section) performs better.
4.2 Natural Language Generation
Although the 4-bit models suffer in SID accuracy, Table 2 shows that the 0.5B 4-bit variant has the highest BERTScore (0.88). This indicates that while it may misclassify device/service fields, the generative text can still be fluent and semantically close to the target. Meanwhile, the 8-bit versions drop in BERTScore for the 0.5B model (0.79) and remain steady for the 1.5B model (0.83). Qualitative samples show that small changes in quantization can shift the style and lexical choices of the generated text.
4.3 Inference Time and Memory
Table 3 summarizes the inference speed across model size, quantization, and CPU core settings. The 8-bit model is only slightly faster than the 16-bit model (5.5 s vs. 6.25 s on 4 cores for the 0.5B). Doubling CPU cores from 2 to 4 reduces latency roughly by half. The 1.5B model takes longer (up to 10–17 s per query), which may be borderline for real-time usage in multi-turn dialogues.
Model Accuracy BERTScore
Qwen2.5-0.5B 84.0 0.68
Qwen2.5-1.5B 86.4 0.66
Table 5: Results Human-Generated Queries. Accuracy and BERTScore over 81 real-user queries.
4.4 Out-of-Domain Intents
In Table 4, we evaluate 60 OOD queries that mention either novel device types or services not appearing in the training set. The 0.5B model scores 80.0% accuracy vs. 86.7% for the 1.5B model, with BERTScores of 0.76 and 0.74 respectively. The results suggest that the 1.5B model generalizes somewhat better to unfamiliar domains, though both degrade compared to in-domain performance.
4.5 Human Prompts
Finally, we tested each model on 81 human-written prompts. Ten participants (ages 23–69) contributed typical commands they would issue to a home assistant, including incomplete or ambiguous phrasing. Table 5 shows that the 0.5B model achieves 84.0% accuracy, whereas the 1.5B model is slightly higher at 86.4%. BERTScores are around 0.66–0.68. The gap vs. synthetic data reflects real-user queries with more variation and noisy data.
5 Discussion
Despite near-perfect performance on the synthetic test set, Table 4 and 5 reveal a drop to 80–86% accuracy in real or out-of-domain queries. This discrepancy likely stems from the difficulty of handling spontaneous human phrasing, missing location or device details, and genuinely novel device types. Still, the results surpass the SVC and DistilBERT baseline.
Interestingly, while the 4-bit model can generate fluent natural language responses (often scoring the highest BERTScore in the 0.5B case), its classification accuracy suffers. This underscores that quantizing a model to extreme levels can degrade structured predictions more than open-ended text generation.
Regarding speed, the 1.5B model yields consistent accuracy gains on OOD data but also increases inference time by up to 2–3$\times$. For single-turn commands, 5–6 seconds per query might be acceptable, but multi-turn dialogue would require faster or more efficient strategies. Future work may explore parameter-efficient fine-tuning, context truncation, or advanced quantization (e.g., 8-bit + partial 4-bit layering) to reduce inference times.
6 Conclusion
We present that LLMs can simultaneously perform SID and natural language response generation for a home automation domain. Experiments on an 8GB RAM, CPU-only environment show that 8-bit quantization largely preserves in-domain accuracy (up to 99%) and strong text fluency, while 4-bit introduces significant classification errors despite retaining good generative capability. We further demonstrate promising generalization to human-written prompts and out-of-domain tasks, with accuracy around 80–86%. However, per-query inference times of 5–6 seconds indicate that LLM-based assistants, as implemented here, are not yet ideal for fast multi-turn dialogues on edge devices. Future work can refine these models for faster, more memory-efficient inference, enabling privacy-preserving yet flexible home automation assistants.
Limitations
Our use of synthetic data may limit the diversity of user prompts; while we partially mitigated this with human-written queries, data coverage remains a challenge. The model also relies on structurally valid JSON output. Real-world usage may need fallback logic to handle malformed or incomplete responses. Moreover, we focus on a single domain (home automation); scaling to broader or open-ended tasks likely requires larger models and may degrade performance under CPU-only constraints.
Ethical Considerations
We do not foresee any major ethical issues with this work. The primary domain is home automation, and the dataset is synthetic or user-provided under informed consent. Nonetheless, deploying generative models in user-facing applications requires caution regarding hallucinated or incorrect responses, as well as user data privacy.
Acknowledgments
MZ and JB are supported by the research grant (VIL57392) from VILLUM FONDEN.
Appendix A Data Examples
We show a data example in Figure 1.
{minted}
[…]
Figure 1: Data Example. In the figure, we show a data example from the acon96 (2024) dataset.
Appendix B Data Distribution Detailed
We show a more detailed distribution of the dataset in Table 6.
Class Total Dataset Test
climate.set_fan_mode 1080 0
climate.set_humidity 1080 0
climate.set_hvac_mode 1080 0
climate.set_temperature 1000 0
cover.close 385 35
cover.open 395 40
cover.stop 320 25
cover.toggle 365 25
fan.decrease_speed 360 60
fan.increase_speed 300 40
fan.toggle 390 85
fan.turn_off 390 70
fan.turn_on 405 60
light.toggle 450 90
light.turn_off 2535 600
light.turn_on 11940 150
lock.lock 200 125
lock.unlock 185 125
media_player.media_next_track 55 25
media_player.media_pause 55 25
media_player.media_play 70 25
media_player.media_previous_track 55 25
media_player.media_stop 55 25
media_player.turn_off 25 25
media_player.turn_on 40 40
media_player.volume_down 65 35
media_player.volume_mute 60 30
media_player.volume_up 85 40
switch.toggle 250 50
switch.turn_off 500 175
switch.turn_on 540 165
timer.cancel 600 0
timer.start 600 0
todo.add_item 1560 0
vacuum.pause 15 0
vacuum.return_to_base 150 0
vacuum.start 370 220
vacuum.stop 15 0
Table 6: Detailed Class Distribution Service. Total Dataset vs. LLM Test Subset